A population of values has a normal distribution with μ=161.9μ=161.9 and σ=55.7σ=55.7. You intend to draw a random sample of size n=61n=61.

What is the mean of the distribution of sample means?
μ¯x=μx¯=  (Enter your answer as a number accurate to 4 decimal places.)

What is the standard deviation of the distribution of sample means?
(Report answer accurate to 4 decimal places.)
σ¯x=σx¯=
Answer #1

Answer:

Given that:

A population of values has a normal distribution with μ=161.9 and σ=55.7

μ=161.9 , σ=55.7 , n= 61

Sampling distribution of sample mean is Vi=

Mi = 161.9

Sampling distribution of sample standard deviation is u/0 = 10

0; = 55.7/161

Oj = 55.7/7.8102

0 = 7.1316